The cost of a long distance phone call is $0.36 for the first min. and $0.21 for each additional min. or portion thereof. Write an inequality representing the number of min. a person could talk without going over $3.00 McGraw Hill, CH. 2. Ex. 7, Problem number 94. Thanks for your help!!!!!!!
Answer by checkley71(8403)

36+21x<=300 or 21x=300-36 or 21x=264 or x=264/21 or x=12.57 therefore a total of 1+12=13 minutes are the max amount of time without exceeding the $3.00 limit.